Solve the following exponential equation. Express irrational solutions in exact form and as a decimal rounded to the three decimal places.
8(2^5x)=9
What is the exact answer?
Select the comment choice below nd if necessary, fill in the answer box to complete your choice.
A. The solution set is?
(simplify answer.Type exact answer)
or
B. There is no solution

Answer by rothauserc(4718)

8 * (2^(5x)) = 9
:
2^3 * (2^(5x)) = 9
:
2^(5x+3) = 9
:
take log base 2 of both sides of =
:
5x + 3 = log(9)/log(2)
:
5x = (log(9)/log(2)) - 3
:
x = (log(9)/(5 * log(2)) - (3/5)
:
*********
A) 0.034 - Real solution
*********
: